### v1

# Unified Manipulability and Compliance Analysis of Modular Soft-Rigid Hybrid Fingers

Abstract
This paper presents a unified framework to analyze the manipulability and compliance of modular soft-rigid hybrid robotic fingers. The approach applies to both hydraulic and pneumatic actuation systems. A Jacobian-based formulation maps actuator inputs to joint and task-space responses. Hydraulic actuators are modeled under incompressible assumptions, while pneumatic actuators are described using nonlinear pressure–volume relations. The framework enables consistent evaluation of manipulability ellipsoids and compliance matrices across actuation modes. We validate the analysis using two representative hands: DexCo (hydraulic) and Edgy-2 (pneumatic). Results highlight actuation-dependent trade-offs in dexterity and passive stiffness. These findings provide insights for structure-aware design and actuator selection in soft-rigid robotic fingers.
keywords:
Soft-rigid Hybrid Actuation, Manipulability Analysis, Compliance Modeling, Modular Robotic Fingers
thanks: Joint first author
, , , and
1 Introduction
Achieving dexterous and robust manipulation in unstructured environments remains a core challenge in robotics Billard and Kragic (2019). As robotic hands advance beyond simple grasping toward versatile in-hand manipulation, there is growing demand for systems that combine precise control, adaptability, compliance, and safe physical interaction. Soft robotic systems excel in adaptive grasping through material compliance and morphology-driven control but face limitations in precision tasks. Soft-rigid hybrid systems, which integrate compliant actuators with rigid linkages, offer a balanced approach—enabling both fine control and structural adaptability.
A variety of soft and soft-rigid robotic hands has emerged over the past decade. Pneumatic systems like the RBO Hand Deimel and Brock (2016) and the Pisa/IIT SoftHand Catalano et al. (2014) enable compliant grasping via soft inflation or tendon-driven underactuation. Hybrid hands enhance control and robustness by combining soft actuators with rigid structures, including origami mechanisms Li et al. (2019), layer jamming Manti et al. (2016), bellow-type encapsulation Lee et al. (2024), and geometric confinement Zhang et al. (2020). Variable stiffness Wolf et al. (2015) and programmable actuation Polygerinos et al. (2017) further boost adaptability.
Our prior work explores both soft and soft-rigid hands. The gripper in Zhou et al. (2017) uses textured dual-chamber fingers for low-pressure adaptation. The BCL-26 hand Zhou et al. (2019) achieves human-like motion with 26 fiber-reinforced soft DOFs. We incorporated stiffness tuning via particle jamming in Zhou et al. (2020). The DexCo Hand Zhou et al. (2024) uses hydraulic actuation for forceful, controllable motion, while the Edgy-2 Hand Zhou et al. (2018) emphasizes modular pneumatic flexibility.
Despite recent advances, many studies remain focused on empirical evaluation and lack generalizable models for analyzing or comparing design variants. While task-specific tuning may suffice for grasping, manipulation tasks demand predictive insights into how structure and actuation affect performance. Two key metrics—manipulability, describing motion capability Doty et al. (1995), and compliance, capturing passive deformation under force Aukes et al. (2014)—are essential for assessing dexterity and adaptability. Though common in rigid and underactuated systems, these metrics are rarely jointly analyzed in soft-rigid hybrid designs, especially across different actuation modes. This gap hinders structure-aware design and cross-system performance generalization.
To address this, we propose a unified analytical framework for evaluating both manipulability and compliance in modular soft-rigid hybrid fingers actuated by hydraulic or pneumatic systems. We formulate a Jacobian-based model that maps actuator inputs to joint and task-space behaviors. Hydraulic systems are modeled using a linear approximation under incompressible flow, while pneumatic actuators are captured using nonlinear pressure–volume relationships derived from ideal gas behavior Polygerinos et al. (2017). This formulation enables consistent computation of manipulability ellipsoids and compliance matrices across actuation types. We validate the proposed framework on two representative systems: the hydraulically actuated DexCo Hand and the pneumatically actuated Edgy-2 Hand. Simulation and experimental results reveal actuation-dependent trade-offs in directional dexterity and passive stiffness, offering insights for structure-aware analysis, actuator selection, and modular finger design in soft-rigid robotic fingers.
Figure 1: (a) Visualization of the model that approximates linear soft actuators. Dashed lines labeled with J1, J2 and J3 marks the positions of the passive revolute joints. Solid lines labeled with L1, L2 and L3 are the linear soft actuator. (b)The top down view of one of the fingers of the Dexco hand when $\theta_{1}=0$. The parameters used in (1) and (2) are labeled.
2 Manipulability Analysis
To analyze the manipulability of the soft-rigid hybrid fingers, we begin with an approximate model that represents the soft actuators as soft linear actuators. Considering the inherent flexibility of soft actuators, we further assume that the soft linear actuators are connected to the knuckles of the DexCo hand via universal joints at both ends. Given the attachment points of the actuators in the local frame of each knuckle, and the poses of the revolute joints, the length of each linear soft actuator corresponds to the distance between its respective attachment points. A visualization of this model is provided in Fig. 1a.
This section analyzes the kinematic mapping between the lengths of the linear soft actuators, denoted as $l_{1}$, $l_{2}$ and $l_{3}$ and the poses of the passive revolute joints, denoted as $\theta_{1}$, $\theta_{2}$, $\theta_{3}$. Based on visual inspection of Fig. 1, this mapping can be decomposed into two parts: $\theta_{3}$ depends solely on $l_{3}$ while $\theta_{1}$ and $\theta_{2}$ are jointly determined by $l_{1}$ and $l_{2}$. The remainder of this section discusses each of these relationships in detail and validates them via hardware experiments.
2.1 Dual Soft Actuation Case
This subsection investigates the kinematic relationship between $l_{1}$, $l_{2}$ and the joint angles $\theta_{1}$, $\theta_{2}$. Through geometric analysis, this relationship can be expressed as follows:
$\begin{split}l^{2}_{1}&=A^{2}+(H\cos{\theta_{1}}-V\sin{\theta_{1}}-S)^{2}\\ &+(H\sin{\theta_{1}}+V\cos{\theta_{1}}-V)^{2}\\ l^{2}_{2}&=A^{2}+(H\cos{\theta_{1}}+V\sin{\theta_{1}}-S)^{2}\\ &+(H\sin{\theta_{1}}-V\cos{\theta_{1}}+V)^{2}\\ \end{split}$ (1)
Here, $A$, $H$, and $V$ represent the axial, horizontal, and vertical offsets, respectively, from joint $J_{1}$ to joint $J_{2}$, as functions of $\theta_{2}$ when $\theta_{1}=0$. These offsets can be computed as follows. The parameter $R$ denotes the radial distance from the attachment points of actuators $L_{1}$ and $L_{2}$ to the axis of joint $J_{2}$. The quantities $D_{1}$ and $D_{2}$ represent the offsets between the actuator attachment surface and the rotation axes. Fig. 1b marks a visualization of these parameters. Lastly, $L$ is the distance between the actuator endpoints on the attachment surface.
$\begin{split}A&=R(1-\cos{\theta_{2}})+D_{2}\sin{\theta_{1}}\\ H&=D_{1}+R\sin{\theta_{2}}+D_{2}\cos{\theta_{2}}\\ V&=\frac{L}{2}\end{split}$ (2)
Figure 2: Surface plot that maps $l1$ and $l2$ into $\theta_{1}$ and $\theta_{2}$. Yellow lines represent sets of $l_{1}$ and $l_{2}$ combinations where $\theta_{1}$ remains the same, while blues lines are sets where $\theta_{2}$ remains the same.
Fig. 2 illustrates the mapping between $(l_{1},l_{2})$ and $(\theta_{1},\theta_{2})$. We fix $\theta_{1}$ at $-15^{\circ}$, $-5^{\circ}$, $5^{\circ}$, and $15^{\circ}$, and plot the corresponding contour lines on the resulting surface. By vertically projecting these contour lines onto the $\theta_{2}$ surface, we obtain curves that represent the achievable range of $\theta_{2}$ through appropriate adjustment of $l_{1}$ and $l_{2}$. Similarly, another set of curves is generated to represent the achievable range of $\theta_{1}$ while holding $\theta_{2}$ fixed. The analysis is constrained to the range $[-20^{\circ},20^{\circ}]$ for both $\theta_{1}$ and $\theta_{2}$, as configurations beyond this range may result in self-collisions. The plots show that for each fixed value of $\theta_{1}$ or $\theta_{2}$, the other joint can still achieve the full range of $[-20^{\circ},20^{\circ}]$ through suitable adjustment of $l_{1}$ and $l_{2}$. This suggests that the pose of joint $J_{1}$ or $J_{2}$ has limited influence on the reachable workspace of the other joint.
Figure 3: (a): The projected curves for contours from A to B and from C to D. The positions of A, B, C, D are marked in Fig. 2. (b): J3 pose against soft actuator L3 length.
To further analyze the relationship between actuator inputs and joint configurations, we plot the projected curves along the contour directions, as shown in Fig. 3a. The two contour paths are defined from point A to B and from point C to D, as indicated in Fig. 2. The x-axis values in these plots correspond to a normalized actuator motion (denoted as $NM$). Taking the contour from A to B as an example, for a given point $(l_{1},l_{2},\theta_{1})$ on the contour, and letting $(l_{1,A},l_{2,A},\theta_{1})$ denote the coordinates at point A, the normalized actuator motion is computed as follows:
$NM=\sqrt{(l_{1}-l_{1,A})^{2}+(l_{2}-l_{2,A})^{2}}$ (3)
As shown in the left side of Fig. 3, the motion of joints $J_{1}$ and $J_{2}$ exhibits an approximately linear relationship with respect to the normalized actuator motion, with a slope of approximately $3.5^{\circ}/\text{mm}$. This slope represents a balanced value, neither too steep nor too shallow, indicating that the soft actuators can effectively drive the passive joints without requiring excessive effort. At the same time, the passive joints are not overly sensitive to small variations in actuator motion, contributing to stable and controllable behavior.
Figure 4: Normalized manipulability distribution.
Fig. 4 presents the Jacobian analysis over the entire workspace. The Jacobian is computed using the following equation, which characterizes the actuator effort required to produce joint motions in arbitrary directions.
$J=\left|\det{\begin{bmatrix}\frac{\partial\theta_{1}}{\partial l_{1}}&\frac{\partial\theta_{1}}{\partial l_{2}}\\ \frac{\partial\theta_{2}}{\partial l_{1}}&\frac{\partial\theta_{2}}{\partial l_{2}}\end{bmatrix}}\right|=\left(\left|\det{\begin{bmatrix}\frac{\partial l_{1}}{\partial\theta_{1}}&\frac{\partial l_{1}}{\partial\theta_{2}}\\ \frac{\partial l_{2}}{\partial\theta_{1}}&\frac{\partial l_{2}}{\partial\theta_{2}}\end{bmatrix}}\right|\right)^{-1}$ (4)
The partial derivatives are directly obtained from (1). To eliminate the influence of unit selection, the Jacobian is normalized by dividing it by its value at the reference configuration where $\theta_{1}=\theta_{2}=0$. As shown in Fig. 4, the normalized Jacobian remains well-conditioned across the workspace, without exhibiting any extreme or anomalous values.
Figure 5: Experimental setup. The joint angles are captured from ArUco code (Romero-Ramirez et al. (2018)).
Figure 6: Linear regressions for hydraulic actuator lengths $l^{e}$ against syringe piston positions $Q^{e}$ in the experiments.
2.2 Single Soft Actuation Case
This section investigates the relationship between $l_{3}$ and $\theta_{3}$. Through direct analytical computation, the following expression can be derived:
$\begin{split}l_{3}^{2}&=H_{3}^{2}+V_{3}^{2}\\ H_{3}&=D_{3}+R\sin{\theta_{3}}+D_{4}\cos{\theta_{3}}\\ V_{3}&=R(1-\cos{\theta_{3}})+D_{4}\sin{\theta_{3}}\end{split}$ (5)
The parameters $D_{3}$ and $D_{4}$ represent the offsets between the soft actuator attachment surface and the rotational axes, as illustrated in Fig. 1b. From this configuration, it can be observed in Fig. 3b that $\theta_{3}$ exhibits an approximately linear relationship with $l_{3}$, with a slope of about $5^{\circ}/\text{mm}$. This suggests that joint $J_{3}$ responds to actuator $L_{3}$ with a balanced sensitivity, neither overly responsive nor excessively resistant to input, indicating effective and stable actuation.
2.3 Experimental Validation with Hydraulic Actuators
To validate the models proposed in (1) and (5), we conducted experiments using a physical DexCo hand equipped with hydraulic actuators $L_{1}$, $L_{2}$, and $L_{3}$. These actuators are filled with incompressible liquid and actuated by three independent external syringes. Let $Q_{1}$, $Q_{2}$, and $Q_{3}$ denote the piston positions of the corresponding syringes. An increase in $Q$ indicates that liquid is being withdrawn from the actuator, resulting in a decrease in actuator length.
Due to the near-incompressibility of the liquid under room temperature and moderate pressure, the actuator lengths $l$ are expected to maintain an affine relationship with the piston positions, i.e., $l=kQ+b$, where $k$ and $b$ are scalar coefficients determined experimentally. Based on this assumption, we collected 200 data samples of the form $(Q_{1}^{e},Q_{2}^{e},Q_{3}^{e},\theta_{1}^{e},\theta_{2}^{e},\theta_{3}^{e})$, and computed the corresponding expected actuator lengths $l_{1}^{e}$, $l_{2}^{e}$, and $l_{3}^{e}$. Linear regression was then performed between the pairs $(Q_{1}^{e},l_{1}^{e})$, $(Q_{2}^{e},l_{2}^{e})$, and $(Q_{3}^{e},l_{3}^{e})$, as illustrated in Fig. 6. The results demonstrate a strong linear correlation between actuator lengths and piston positions, with high coefficients of determination, validating the geometric models in (1) and (5).
2.4 Potential Extension to Pneumatic Actuators
It is worth noting that the geometric modeling presented in Sections 2.1 and 2.2 constitutes a generalizable framework that can also be applied to pneumatic actuators. While air is typically treated as a compressible fluid, it is common practice to establish a mapping from the piston position $Q$ and the pressure within the syringe $P$ to the actuator length $l$ through experimental calibration. Once this mapping is identified, it becomes possible to derive the relationship between $(Q,P)$ and the resulting joint configurations.
Figure 7: Illustration of compliance modeling during the force estimation stage in actuation space.
3 Compliance Modeling
While manipulability characterizes the motion capability of the fingers, compliance captures how they respond to external forces—an equally critical factor in soft-rigid design. Due to the inherent compliance of soft materials, both hydraulic and pneumatic actuators exhibit compliant behavior during interaction. In this section, we focus on the static analysis and compliance modeling of the DexCo hand, mapping from the actuation space ($\boldsymbol{a}$) to the configuration space ($\boldsymbol{\tau}$). We develop models to characterize the compliance introduced by both hydraulic and pneumatic actuation, and validate these models through simulations involving object manipulation.
3.1 Hydraulic Compliance Modeling
In hydraulic systems, compliance arises primarily from the compressibility of the working fluid, elasticity of the fluidic chamber and tubing, and mechanical compliance of the structure. Although hydraulic actuation is often regarded as high-stiffness and low-compliance, for soft-rigid hybrid systems, even small compliance effects can significantly influence control and force interaction performance.
The hydraulic force in soft actuators arises from the elastic deformation of the material in the radial direction during compression or extension, similar to inflating a balloon (Fig. 7).
We assume under an initial height $h_{0}$ of the soft actuator, the volume of water remains constant during motion, i.e., $V=V_{0}\triangleq\text{constant}$. Based on this assumption, the geometric constraint is given by:
$V_{0}$ $=\frac{\pi h_{0}}{3}\left[\left(\frac{D_{0}}{2}\right)^{2}+\left(\frac{D_{0}}{2}\right)^{2}+\frac{D_{0}d_{0}}{4}\right]$ (6)
$V$ $=\frac{\pi h}{3}\left[\left(\frac{D}{2}\right)^{2}+\left(\frac{D}{2}\right)^{2}+\frac{Dd}{4}\right]$
$V$ $=V_{0}$
To evaluate the radial deformation, we apply thin-walled pressure vessel theory. As illustrated in Fig. 7, the hoop stress $\sigma_{1}$ is calculated along the actuator’s axial profile:
$\sigma_{1}=\frac{Pr}{t}$ (7)
where $r$ is the local radius at height $y$, $P$ is the internal pressure, and $t$ is the wall thickness. According to the generalized Hooke’s law, the circumferential strain $\varepsilon_{1}$ is:
$\varepsilon_{1}=\frac{\sigma_{1}}{E}=\frac{\Delta_{1}}{2\pi r}$ (8)
where $\Delta_{1}$ is the change in circumference due to pressure. Using (7) and (8), we derive the expressions for the deformed inner and outer diameters under pressure $P$:
$d$ $=d_{0}+\frac{2P}{Et}\left(\frac{d_{0}}{2}\right)^{2}$ (9)
$D$ $=D_{0}+\frac{2P}{Et}\left(\frac{D_{0}}{2}\right)^{2}$
where the internal pressure is defined as $P\triangleq-\frac{4}{\pi d_{0}^{2}}F_{h}$. Substituting (9) into (6), we can solve for the hydraulic force $F_{h}=F_{h}(h_{0},\Delta y)$ by solving a resulting quadratic equation.
To characterize the pressure-to-displacement relationship of the hydraulic actuator, we define two compliance-related parameters. The first is the hydraulic compliance $C_{h}$, which reflects the actuator’s deformation response to external force and is expressed as displacement per unit force. To translate this into a form compatible with pressure input modeling, we introduce the effective linear compliance $C_{l}=C_{h}/A$, where $A$ denotes the actuator’s effective cross-sectional area. This conversion yields a linear input-output relationship between internal pressure $P$ and output displacement $l$, given by $l=C_{l}\cdot P$. This formulation allows for consistent integration of actuator compliance into the Jacobian-based framework for joint-space stiffness and task-space compliance analysis.
In joint space, the Jacobian matrix $J\in\mathbb{R}^{n\times m}$ relates joint displacements $\Delta q\in\mathbb{R}^{m}$ to end-effector displacements $\Delta l\in\mathbb{R}^{n}$ via $\Delta l=J\cdot\Delta q$, as defined in Equation (4). Under an external force $\Delta f\in\mathbb{R}^{n}$, the resulting torque deviation in joint space $\Delta\tau\in\mathbb{R}^{m}$ can be expressed as:
$\Delta\tau=J^{T}\cdot\Delta f=J^{T}\cdot\frac{1}{C_{l}}\cdot\Delta l=\frac{1}{C_{l}}J^{T}J\cdot\Delta q$ (10)
Here, $C_{l}$ is the effective linear compliance that maps input pressure to displacement. Based on this relationship, we derive the joint-space stiffness matrix $K_{q}\in\mathbb{R}^{m\times m}$ and its inverse (i.e., joint-space compliance matrix $C_{q}$) as:
$K_{q}=\frac{1}{C_{l}}J^{T}J,\quad C_{q}=K_{q}^{-1}=C_{l}\cdot(J^{T}J)^{-1}$ (11)
To analyze system behavior in task space, we transform these joint-space matrices using a task-space Jacobian $J_{a}\in\mathbb{R}^{n\times m}$, leading to:
$\begin{gathered}K_{a}=J_{a}^{T}K_{q}J_{a}\\ J_{a}C_{a}J_{a}^{T}=C_{q}\end{gathered}$ (12)
where $K_{a}$ and $C_{a}$ are the task-space stiffness and compliance matrices, respectively. This transformation reveals how global actuator compliance $C_{l}$ interacts with configuration-dependent geometry (via $J$ and $J_{a}$) to shape the directional compliance profile in task space.
This result is analogous to classical manipulability-based compliance analysis, revealing how the hydraulic compliance $C_{l}$ globally scales the configuration-dependent compliance profile.
3.2 Pneumatic Actuation Model
In pneumatic systems, the input pressure is indirectly regulated via flow control, and the resulting output force is strongly dependent on chamber volume, which changes with deformation. Therefore the geometric constrain assumption made in hydraulic actuation is not ever applicable (6).
Assuming quasi-static behavior and ideal gas behavior, we model the internal pressure as:
$P(l)=\frac{nRT}{V(l)}=\frac{nRT}{V_{0}+\alpha l}$ (13)
where $l$ is the deformation-dependent variable (e.g., finger length or curvature), and $\alpha$ is a volume gain coefficient. The corresponding output force is:
$f(l)=\beta P(l)=\beta\cdot\frac{nRT}{V_{0}+\alpha l}$ (14)
This nonlinear relation suggests a decreasing output force with increasing deformation, introducing variable stiffness properties in the actuation. When mapped to joint space via the Jacobian, the torque is:
$\tau=J^{T}(q)f(q)$ (15)
highlighting the pressure-volume-deformation coupling in the control of soft-rigid pneumatic systems.
3.3 Simulate Compliance Motion
In simulation, we focus on simulating the hydraulic compliance. The variable stiffness property inherent in the DexCo hand can be modeled either as a linear variable elastic force, expressed as follows:
$\tau_{k}=-k(\theta_{0})\,\Delta\theta$ (16)
where $\theta_{0}$ is the joint angle in the absence of external forces, $\theta$ is the current joint angle, and $\Delta\theta=\theta-\theta_{0}$ denotes the angular deformation. The linear elastic force in Eq. 16 can be transformed into Cartesian space using the stiffness matrix $K_{a}=J_{a}^{\top}K_{q}J_{a}$, where $J_{a}$ is the Jacobian and $K_{q}$ is the joint stiffness matrix.
When modeling the stiffness as a torque input in simulation, it affects the system dynamics as follows:
$M(\theta)\,\ddot{\theta}+H(\dot{\theta},\theta)=\tau_{k}$ (17)
In Eq. 17, the left-hand side represents the inertial, Coriolis, and gravitational forces, which are computed by the physics simulator (gazebo). The right-hand side, $\tau_{k}$, denotes the torque induced by elastic deformation. By incorporating $\tau_{k}$ into the simulation, we are able to replicate the effects of nonlinear, variable stiffness elastic forces on the hand’s behavior, as shown in the Fig. 8B.
In the implemented simulation, we conducted a motion sequence consisting of two phases. First, a compliant caging grasp was performed to securely enclose the object. Then, by increasing the actuation force at the base joint of the left finger while altering the orientation of the fingertip, the object was deliberately pushed away. Figure 8A illustrates the commanded joint angles alongside the actual joint responses. Due to the variable compliance of the system, discrepancies arise between the desired and actual joint angles. Joints 2 and 3 correspond to the universal joint at the finger base, while Joint 4 represents the fingertip joint.
Figure 8: Simulation in gazebo of the compliant motions of the DexCo hand.
4 Conclusion
This paper proposed a unified framework for analyzing the manipulability and compliance of modular soft-rigid hybrid fingers actuated by hydraulic and pneumatic systems. A generalized Jacobian-based formulation was introduced to model the actuator-to-task-space relationship, enabling consistent evaluation of directional dexterity and passive compliance across actuation types. Simulations and experiments on the DexCo and Edgy-2 fingers validated the framework and highlighted key trade-offs between actuation modality, structural configuration, and control sensitivity.
The results provide a foundation for structure-aware and actuation-aware design of soft-rigid fingers, facilitating informed selection and optimization in modular robotic systems. Future work will extend this framework to include dynamic effects such as valve delay and flow regulation in pneumatic actuation, as well as integrate contact modeling for task-level compliance control and grasp planning.